OpenStudy (anonymous):

what is 4/5 of 10/13?

OpenStudy (johnweldon1993):

When you see "of"...multiply \[\large \frac{4}{5} \times \frac{10}{13} = ?\]

OpenStudy (anonymous):

Oh, okay. Thanks John

OpenStudy (johnweldon1993):

No problem!